Need all parts answered step by step.
over time -4 Suppose that five years ago Cisco Systems sold a 15-year bond issue that had a Bond Valuati ct to receive tween the rms deci- of its $1,000 par value and a 7 percent coupon rate. Interest is paid semiannually a. If the going interest rate has risen to 10 percent, at what price would the bonds be selling today? Suppose that the interest rate remained at 10 percent for the next 10 years. What would happen to the price of Ciscos bonds over time? b. ant to Preferred S
Answer #1

A:

Using financial calculator

Input: FV=1000, N = 15-5=10, PMT=7%*1000 = 70 , I/Y = 10%

Find PV as $815.66

B:

After 10 years, the Price of the bond will be its Par value which is $1000. This is the value which will be received at maturity.